### v1

# Inference-friendly Graph Compression for Graph Neural Networks

Abstract.
Graph Neural Networks ($\mathsf{GNNs}$) have demonstrated promising performance in graph analysis. Nevertheless, the inference process of $\mathsf{GNNs}$ remains costly, hindering their applications for large graphs. This paper proposes inference-friendly graph compression ($\mathsf{IFGC}$), a graph compression scheme to accelerate $\mathsf{GNNs}$ inference. Given a graph $G$ and a $\mathsf{GNN}$ $M$, an $\mathsf{IFGC}$ computes a small compressed graph $G_{c}$, to best preserve the inference results of $M$ over $G$, such that the result can be directly inferred by accessing $G_{c}$ with no or little decompression cost. (1) We characterize $\mathsf{IFGC}$ with a class of inference equivalence relation. The relation captures the node pairs in $G$ that are not distinguishable for $\mathsf{GNN}$ inference. (2) We introduce three practical specifications of $\mathsf{IFGC}$ for representative $\mathsf{GNNs}$: structural preserving compression ($\mathsf{SPGC}$), which computes $G_{c}$ that can be directly processed by $\mathsf{GNN}$ inference without decompression; $(\alpha,r)$-compression, that allows for a configurable trade-off between compression ratio and inference quality, and anchored compression that preserves inference results for specific nodes of interest. For each scheme, we introduce compression and inference algorithms with guarantees of efficiency and quality of the inferred results. We conduct extensive experiments on diverse sets of large-scale graphs, which verifies the effectiveness and efficiency of our graph compression approaches.
1. Introduction
[…]
Despite their promising performances, $\mathsf{GNNs}$ incur expensive inference process when $G$ is large  (Chen et al., 2020; Zeng et al., 2021; Zhou et al., 2021). The emerging need for large-scale testing, fine-tuning and benchmarking of graph learning models, require fast inferences of $\mathsf{GNNs}$ under various configurations. Consider the following scenarios.
(1) Inference in large networks. For a graph $G$ with $|V|$ nodes and $|E|$ edges (where $V$ and $E$ refers to its node and edge set), with on average $F$ features per node, an $L$-layered $\mathsf{GNN}$ $M$ may typically take $O(L|E|dF^{2}+L|V|F^{2})$ time  (Chen et al., 2020) (as summarized in Table 3). This can be prohibitively expensive for large real-world graphs.
(2) Real-time Inference. $\mathsf{GNN}$ models have been developed for e.g., traffic analysis (Peng et al., 2020), social recommendation (Fan et al., 2019), energy forecasting (Karimi et al., 2021; Fan et al., 2023; Ahmed et al., 2022), cybersecurity (Zhou et al., 2019), computer vision (Shi and Rajkumar, 2020), and edge devices (Zhou et al., 2021). Such scenarios often require real-time response at e.g., milliseconds  (Zhou et al., 2021). In such cases, even a linear time inference of “small” $\mathsf{GNNs}$ (when $F$ and $L$ are small constants) may still not be feasible for large graphs (when $|V|$ and $|E|$ are large).
(3) Fine-tuning & Benchmarking. Fine-tuning and testing pre-trained $\mathsf{GNNs}$ to adapt them for various domain-specific tasks is a routine process in $\mathsf{GNN}$-based data analysis for e.g., materials sciences, biomedicine, social science, and geosciences (Han et al., 2021; Zhili et al., 2024; Zhou et al., 2020; Reiser et al., 2022; Veličković, 2023). Inference tests of large pool of “candidate” $\mathsf{GNNs}$ over domain-specific graph data (such as knowledge graphs) is a cornerstone in such context. Fast $\mathsf{GNN}$ inference can accelerate large-scale domain-specific testing and benchmarking.
[…]
(a) Social role classification: a hospital network ($G_{1}$) can be “compressed” by merging nodes with equivalent social roles for testing a $\mathsf{GCN}$-based classifier (adapted from (Borgatti and Everett, 1992), node attributes include Age, YoE (years of experience), and Dpt (department)).
(b) Visualization of a fraction of real-world citation network (Hu et al., 2020) ($|G|=1,335,586$) and its compressed counterpart ($|G_{c}|=148,887$) for a $\mathsf{GraphSAGE}$-based node classification. $88.6\%$ of nodes and edges are compressed, reducing inference cost by $92.0\%$, achieving $12.5$ times speed-up with up to $6.7\%$ loss of accuracy.
Figure 1. Compression Scheme to scale node classification.
Compressing graphs for $\mathsf{GNN}$ inference. Unlike prior “model-specific” approaches, we propose a model-agnostic, “once-for-all” graph compression scheme to accelerate $\mathsf{GNN}$ inference, for a set of $\mathsf{GNNs}$. Consider a set of $\mathsf{GNNs}$ $\mathbb{M}$ (a $\mathsf{GNN}$ “class”) with the same form of inference, which apply the same “type” of the node update function but only differs in model weights (see Example 1). The inference of a $\mathsf{GNN}$ $M$ over $G$ can be characterized as an “inference query” (Barceló et al., 2020; Geerts, 2023), which invokes the inference of ${\mathcal{M}}$ to compute ${\mathcal{M}}(G)$.
We advocate an “inference-friendly” graph compression scheme for $\mathsf{GNN}$ inference at scale. Given $\mathbb{M}$ and a large graph $G$,
  • $\circ$
    It uses a compression function ${\mathcal{C}}$ to compute a smaller counterpart $G_{c}$ of $G$ “once-for-all”, for any ${\mathsf{GNN}}{\mathcal{M}}\in\mathbb{M}$;
  • $\circ$
    For any inference query that requests ${\mathcal{M}}(G)$ for a specific $\mathsf{GNN}$ ${\mathcal{M}}\in\mathbb{M}$, it performs an inference directly over $G_{c}$ instead of $G$ to compute ${\mathcal{M}}(G_{c})$, with a reduced time cost, such that ${\mathcal{M}}(G_{c})$ (approximately) equals ${\mathcal{M}}(G)$.
Such a compression is desirable: (1) It readily reduces the cost for any single inference query that computes ${\mathcal{M}}(G)$; (2) An inference query often does not require the entire output ${\mathcal{M}}(G)$ but only a fraction ${\mathcal{M}}(G,V_{T})\subseteq{\mathcal{M}}(G)$ of a specified test (node) set $V_{T}$ of interests; (3) Multiple inference queries can be posed to request output from different $\mathsf{GNNs}$ in $\mathbb{M}$ in $G$. For any workload with inference queries that specify any ${\mathsf{GNN}}$ ${\mathcal{M}}\in\mathbb{M}$ and any $V_{T}$ from $G$, one only need to compute $G_{c}$ once, to reduce the total inference cost of the workload. These benefits applications in large-scale tests over large graphs, real-time inference and benchmarking, as aforementioned.
While desirable, is such a compression scheme doable? We illustrate a case in the following example.
Example 1.
[…]
We take a closer look at the update function $M_{v}$ at layer $k$:
$X_{v}^{k}=\sigma(\Theta\cdot\sum_{u\in N(v)}X_{u}^{k-1})$
where $X_{v}^{k-1}$ (resp. $X_{v}^{k}$) is the embedding of a node $v$ at the $(k-1)$-th (resp. $k$-th) layer; $\sigma$ is an activation function, $N(v)$ refers to the neighbors of node $v$, and $\Theta$ refers to the learned weight matrix (same across all layers in the $\mathsf{GNN}$).
[…]
Note that feature equivalence does not necessarily mean that $x_{1}$ and $x_{2}$ have exactly the same attribute values as input. For example, while $n_{1}$ and $n_{2}$ refer to a $26$ years old and a $28$ years old nurse, respectively, their ages, years of experience (YoE), and department (Dpt) fall in the same group via (categorical or one-hot) encoding. Hence, they have the same input feature.
By “recursively” merging all such node pairs into a “group node” that are connected to neighbors that are also indistinguishable groups (e.g., $[s]=\{s_{1},s_{2}\}$, $[d]=\{d_{1},d_{2}\}$, $[n]=\{n_{1},n_{2}\}$, and $[v]=\{v_{1},v_{2}\}$), a smaller graph $G_{c}$ can be obtained. An inference directly over $G_{c}$ can yield the same output for the test nodes $v_{1}$ and $v_{2}$ for ${\mathcal{M}}$ without decompression. To see this, one just need to “recover” their original value with a constant factor $2$ (their original degrees) as auxiliary information at query time, for each layer:
$X_{v_{1}}^{k}=X_{v_{2}}^{k}=2\times X_{[v]}^{k}=2\times\sigma(\Theta\cdot X_{[n]}^{k-1})$
Such aggregated neighborhood information (e.g., degrees, edge weights/attentions, or hyper-parameters) can be readily “remembered” at compression time, and be retrieved in constant time. This indicates an overall cheaper inference cost, and an exact query-time restore of the original embedding, for any node in $G$.
[…]
The above example verifies the possibility of a graph compression scheme by finding and merging node pairs that are indistinguishable for the inference process of $\mathsf{GNNs}$. Our study verifies that real-world graphs are indeed highly compressible with such structures, and if compressed, well preserve inference output with no or small loss of accuracy, and meanwhile significantly reduce unnecessary inference computation (see Fig. 1 (b)).
Contributions. Our main contributions are as follows.
(1) We formally introduce inference-friendly graph compression scheme ($\mathsf{IFGC}$), a as a general scheme to scale $\mathsf{GNN}$ inference to large graphs. We characterize $\mathsf{IFGC}$ with an inference equivalence relation, which captures the nodes with embeddings that are indistinguishable for the inference process using the same type of node update function. We then introduce a sufficient condition for the existence of $\mathsf{IFGC}$, which specifies $G_{c}$ as the quotient graph of $G$ induced by the inference equivalence relation.
(2) We specify $\mathsf{IFGC}$ for representative $\mathsf{GNNs}$ classes. We first introduce structural preserving compression ($\mathsf{SPGC}$), that enforces node embedding equivalence and neighborhood connectivity. We show it computes a compressed graph $G_{c}$ in $O(|E|\log|V|)$ time, which can be directly processed by the inference process to retrieve the original results without decompression. We further justify $\mathsf{SPGC}$ by showing that it can produce a unique, smallest $G_{c}$ up to graph isomorphism among compressed graphs. We also show that $\mathsf{SPGC}$ preserves the discriminability of the $\mathsf{GNNs}$.
(3) We further introduce two configurable variants of $\mathsf{IFGC}$ to allow flexible trade-off between compression ratio and the quality of inference output. (a) The $(\alpha,r)$-$\mathsf{SPGC}$ groups nodes with similar features (determined by a threshold $\alpha$), that also have similar counterparts within their $r$-hop neighbors. (b) The anchored-$\mathsf{SPGC}$ ($\mathsf{ASPGC}$) adapts $\mathsf{SPGC}$ to an “anchored” node set of user’s interests and preserve the inference results for such nodes only rather than the entire node set. For both variants, we introduce efficient compression and inference algorithms.
(4) We experimentally verify the effectiveness and efficiency of our graph compression schemes. We show that with cheap “once-for-all” compression, our compression methods can significantly reduce the inference cost of representative $\mathsf{GNNs}$ such as $\mathsf{GCNs}$, $\mathsf{GAT}$ and $\mathsf{GraphSAGE}$ by 55%-85%, with little to no sacrifice of their accuracy.
Methods Category LB MA IFGS Compression Cost
Dspar (Liu et al., 2023) S $\times$ $\checkmark$ $\checkmark$ $O(\frac{|V|log|V|}{\epsilon^{2}})$
AdaptiveGCN (Li et al., 2021) S $\checkmark$ $\times$ $\times$ $n/a$
NeuralSparse (Zheng et al., 2020) S $\checkmark$ $\checkmark$ $\times$ $O(q|E|)$
SCAL (Huang et al., 2021) C $\checkmark$ $\checkmark$ $\times$ $n/a$
FGC (Kumar et al., 2023) C $\checkmark$ $\checkmark$ $\times$ $O(|V|^{2}|V_{c}|)$
$\mathsf{SPGC}$ (Ours) C $\times$ $\checkmark$ $\checkmark$ $O(|V|+|E|)$
$(\alpha,r)$-$\mathsf{SPGC}$ (Ours) C $\times$ $\checkmark$ $\checkmark$ $O(|V||N_{r}|+|E|)$
$\mathsf{ASPGC}$ (Ours) C $\times$ $\checkmark$ $\checkmark$ $O(|G_{L}|)$
Table 1. Graph compression to accelerate $\mathsf{GNN}$ inference. S: Sparsification, C: Coarsening, LB: Learning-Based, MA: Model-Agnostic, IP: Inference-friendly w. guarantee, $\epsilon$: a constant that controls approximation error, $q$: ##\## of visits of the neighbors per node. $N_{r}$: the largest $r$-hop neighbor set for a node in $G$. $G_{L}$: the subgraph of $G$ induced by $L$-hop of anchored node set $V_{A}$ for $L$-layered $\mathsf{GNNs}$. .
Related work. Several approaches have been developed to accelerate $\mathsf{GNN}$ inference in large graphs (Gao et al., 2022; Zhou et al., 2021; Chen et al., 2021; Peng et al., 2022). Closer to our approach is graph reduction, which simplifies graphs at a small sacrifice of model accuracy (Liu et al., 2022; Hashemi et al., 2024). There are three strategies.
Graph Sparsification. These methods (learn to) remove task-irrelevant edges from input graphs, such that the remaining part preserves the performance of $\mathsf{GNNs}$. For example, AdaptiveGCN (Li et al., 2021) learns an edge predictor to determine and remove task-irrelevant edges to accelerate GNN inference on CPU/GPU clusters. NeuralSparse (Zheng et al., 2020) learns supervised DNNs to remove task-irrelevant edges.  (Chen et al., 2021) proposed a framework to incorporates both model optimization and graph sparsification, which leverages lottery ticket hypothesisto identify subnetworks that can perform as well as the full network. Dspar (Liu et al., 2023) induces smaller subgraphs by removing edges that have similar “importance” (quantified by approximating a resistance measure as in circuits) to preserve graph spectrum.
Graph Coarsening. These methods group and amalgamate nodes into groups, without removing nodes. For example, SCAL (Huang et al., 2021) proposed the use of off-the-shelf coarsening methods LV(Loukas and Vandergheynst, 2018) for scaling up GNN training and theoretically proved that coarsening can be considered a type of regularization and may improve the generalization as well as reduce the number of nodes by up to a factor of ten without causing a noticeable downgrade in classification accuracy. (Kumar et al., 2023) introduced an optimization-based framework (FGC) that incorporates graph matrix and node features to jointly learn a coarsened graph while preserving desired properties such as spectral similarity (Loukas and Vandergheynst, 2018). GRAPE (Xu et al., 2021) is a $\mathsf{GNN}$ variant enhanced with sampled subgraph features from ego networks of automorphic equivalent nodes. It has a different goal of improving accuracy rather than reducing inference costs.  (Bollen et al., 2023) compresses graphs to accelerate $\mathsf{GNN}$ learning, using color refinement (with a case of bisimulation) to merge nodes within bounded radius. Node groups are iteratively refined based on a label encoding that concatenate node label and neighboring group colors. This is similar with $\mathsf{SPGC}$. Nevertheless, no inference algorithm is provided. We show that our $(\alpha,r)$-compression subsumes bisimulation compression.
Graph Sketch. These methods reduce the redundancy in the original graph, generating a skeleton graph that retains essential structural information. For instance, Graph-Skeleton (Cao et al., 2024) constructs a compact, synthetic, and highly-informative graph for the target nodes classification by eliminating redundant information in the background nodes. While this approach effectively reduces the memory usage for a pre-defined set of target nodes, it lacks flexibility and generalizability, as it must be tailored to specific nodes. In contrast, our $\mathsf{SPGC}$ performs one-time compression that applies universally to any set of test nodes $V_{T}\subset V$. NeutronSketch (Liu et al., 2025) focuses on eliminating the redundant information from the training portion of the graph, yet its compression is restricted to the training phase rather than accelerating inference, and does not ensure inference equivalence for the compressed graphs.
Our work differs from existing graph reduction approaches (summarized in Table. 1) in the following. (1) Our methods are model-agnostic and apply to any $\mathsf{GNN}$ that adopt the same inference process, without requiring model parameters, and incur no learning overhead. (2) We specify $\mathsf{IFGC}$ with variants that (approximately) preserve inferred results with invariant properties such as uniqueness and minimality, as well as fast compression and inference algorithms. These are not discussed in prior work. On the other hand, we remark that our scheme can be applied orthogonally: One can readily apply these approaches over compressed graphs from our method to further improve $\mathsf{GNN}$ training and inference. Our proposed $\mathsf{SPGC}$ can be potentially extended to the following learning settings: (1) transductive-learning: training on a compressed graph $G_{c}$ while utilizing our proposed memoization structure ${\mathcal{T}}$ to enable accuracy-preserving inference; and (2) inductive-learning: applying semi-supervised or unsupervised learning techniques to refine the graph structure of $G_{c}$ for a specific task.
2. Graphs and Graph Neural Networks
Graphs. A directed graph $G=(V,E)$ has a set of nodes $V$ and a set of edges $E\subseteq V\times V$. Each node $v$ carries a tuple $T(v)$ of attributes and their values. The size of $G$, denoted as $|G|$, refers to the total number of its nodes and edges, i.e., $|G|$ = $|V|+|E|$.
Graph Neural Networks. A graph neural network (${\mathsf{GNN}}$) ${\mathcal{M}}$ is a mapping that takes as input a featurized representation $G$ = $(X,A)$ to an output embedding matrix $Z$, i.e., ${\mathcal{M}}(G)$ = $Z$. Here $X$ is a matrix of node features, and $A$ is a (normalized) adjacency matrix of $G$ [^1] A feature vector $X_{v}$ of a node $v$ can be a word embedding or one-hot encoding (Gardner et al., 2018) of $T(v)$. $A$ is often normalized as $\hat{A}$ = $A$ + $I$, where $I$ is the identity matrix.
Inference. We take a query language perspective  (Barceló et al., 2020; Geerts, 2023) to characterize the inference process of $\mathsf{GNNs}$. A $\mathsf{GNN}$ inference process is specified as a composition of node update functions.
[…]
$x^{k}_{v}=M_{v}^{k}(\Theta^{k},{\mathsf{AGG}}(X^{k-1}_{u},x^{k-1}_{v},\forall u\in N^{k}(v)))\vspace{-1ex}$
which is specified by (1) the learned model parameters $\Theta^{k}$, (2) an aggregation function ${\mathsf{AGG}}$ (e.g., $\sum$, $\mathsf{CONCAT}$), and (3) the neighbors of $v$ that participate in the inference computation at the $k$-th layer (denoted as $N^{k}(v)$). When $k$=1, $X^{0}_{v}$=$X_{v}\in X$, i.e., the input features.
[…]
We consider $\mathsf{GNNs}$ with fixed, deterministic inference processes. In practice, such $\mathsf{GNNs}$ are desired for consistent and robust performance. For simplicity, we assume that $M_{v}$ specifies a proper set of neighbors that participate the inference process as $N(v)\subseteq\{u|(u,v)\in E$ or $(v,u)\in E\}$. This allows us to include $\mathsf{GNNs}$ that exploits neighborhood sampling (such as $\mathsf{GraphSAGE}$), and directed message passing into discussion. In general, inferences of representative $\mathsf{GNNs}$ are in $\mathsf{PTIME}$ (Chen et al., 2020; Zhou et al., 2021) (see Table 3).
Classes of $\mathsf{GNNs}$. We say a set of fixed, deterministic $\mathsf{GNNs}$ $\mathbb{M}$ belongs to a class of $\mathsf{GNNs}$ $\mathbb{M}^{L}$, if for every $\mathsf{GNN}$ ${\mathcal{M}}\in\mathbb{M}$, (1) ${\mathcal{M}}$ has $L$ layers, and (2) ${\mathcal{M}}$ uses the same form of node update function $M^{k}_{v}$, for each node $v\in V$ and $k\in[1,L]$.
Table 3 summarizes several node update functions in their general forms for mainstream $\mathsf{GNN}$ classes. For example, Graph Convolution Networks ($\mathsf{GCNs}$(Kipf and Welling, 2016) adopt a node update function as $X_{v}^{k}=\sigma(\Theta^{k}(\sum_{u\in\mathcal{N}(v)}\frac{1}{\sqrt{d_{u}d_{v}}}x_{u}^{k-1}))$. Here $d_{u}$ or $d_{v}$ denotes the degree of node $u$ or $v$. $\sigma(.)$ is the non-linear activation function. A class of $\mathsf{GNNs}$ ${\mathsf{GCN}}^{3}$ contains 3-layered $\mathsf{GCNs}$ that adopt such node update function. Note that two $\mathsf{GNNs}$ in the same class can have different $\Theta$ and output, given the same input.
Notation Description
$G=(X,A)$ graph $G$, $X$: feature matrix,
$A$: adjacency matrix
$|G|$ size of $G$; $|G|=|V|+|E|$
${\mathcal{M}}$ a $\mathsf{GNNs}$ model
${\mathcal{M}}(G)$ (resp. ${\mathcal{M}}(G,V_{T})$) output of ${\mathcal{M}}$ over $G$ (resp. test set $V_{T}$)
${\mathcal{C}},{\mathcal{P}}$ compression, post-processing function
$M_{v}$ node update function
$x^{k}_{v}$ embedding of node $v$ at layer $k$
$G_{c}$ compressed graph of $G$
$\alpha,r$ similarity threshold, # hops
$V_{T},V_{A}$ test node set, anchored nodes
$R^{S},R^{(\alpha,r)},R^{A}_{L}$ structural equivalence, $(\alpha,r)$ relation
& anchored relation
Table 2. Summary of Notations.
$\mathsf{GNNs}$ Classes Node Update Function (general form) Training Cost Inference Cost
Vanilla (Scarselli et al., 2008) $X_{v}^{k}=\sigma(\Theta\cdot\text{AGG}(X_{u}^{k-1},\forall u\in\mathcal{N}(v)))$ $O(L|E|+L|V|)$ $O(L|E|+L|V|)$
$\mathsf{GCN}$ (Kipf and Welling, 2016; Chen et al., 2020) $X_{v}^{k}=\sigma(\Theta^{k}(\sum_{u\in\mathcal{N}(v)}\frac{1}{\sqrt{d_{u}d_{v}}}x_{u}^{k-1}))$ $O(L|E|F+L|V|F^{2})$ $O(L|E|F+L|V|F^{2})$
$\mathsf{GAT}$ (Velickovic et al., 2017; Besta and Hoefler, 2024) $X_{v}^{k}=\sigma(\sum_{u\in\mathcal{N}(v)}\alpha_{uv}\Theta^{k}X_{u}^{k-1}))$ $O(L|E|dF^{2}+L|V|F^{2})$ $O(L|E|dF^{2}+L|V|F^{2})$
$\mathsf{GraphSAGE}$ (Hamilton et al., 2017; Chen et al., 2020) $X_{v}^{k}=\sigma(\Theta^{k}\cdot(X_{v}^{k-1}||\text{AGG}(X_{u}^{k-1},\forall u\in\mathcal{N}(v))))$ $O(L|V|dF+L|E|dF^{2})$ $O(L|V|dF+L|E|dF^{2})$
$\mathsf{GIN}$ (Xu et al., 2019; Besta and Hoefler, 2024) $X_{v}^{k}=\sigma(\text{MLP}((1+\gamma)x_{v}^{k-1}+\sum_{u\in\mathcal{N}(v)}x_{u}^{k-1})$ $O(L|E|F+L|V|F^{2})$ $O(L|E|F+L|V|F^{2})$
Table 3. Comparison of Representative $\mathsf{GNNs}$ with node update functions, training cost, and inference cost. $\sigma$: an activation function e.g., ReLU or LeakyReLU. AGG: aggregation function; can be e.g., sum ($\sum$), average (${\mathsf{Avg}}$), or concatenation (||||| |). $L$, $|E|$, $|V|$, $F$, and $d$ denote the number of layers, edges, nodes, features per node, and maximum node degree of $G$, respectively.
3. Inference-friendly Compression
Given a graph $G$ = $(V,E)$, a compressed graph of $G$, denoted as $G_{c}$ = $(V_{c},E_{c})$, is a graph where (1) each node $[v]\in V_{c}$ is a nonempty subset of $V$, and $V$ = $\bigcup_{[v]\in V_{c}}[v]$; and (2) there is an edge $([v],[v^{\prime}])\in E_{c}$, if there are at least one node $v\in[v]$ and $v^{\prime}\in[v^{\prime}]$, such that $(v,v^{\prime})\in E$.
Note that $|V_{c}|\leq|V|$ and $|E_{c}|\leq|E|$. Hence, $|G_{c}|\leq|G|$.
Inference-friendly Graph Compression. Given a set of $\mathsf{GNNs}$ $\mathbb{M}$ and a graph $G$, an inference-friendly graph compression, denoted as $\mathsf{IFGC}$, is a pair $({\mathcal{C}},{\mathcal{P}})$ where
  • $\circ$
    ${\mathcal{C}}$ is a compression function that computes a compressed graph $G_{c}$ of $G$ ($G_{c}$ = ${\mathcal{C}}(G)$);
  • $\circ$
    ${\mathcal{P}}$ is a function that restore the auxiliary information of nodes in $G_{c}$ to their counterparts in $G$; and moreover,
  • $\circ$
    ${\mathcal{M}}(G)$ = ${\mathcal{M}}({\mathcal{P}}(G_{c}))$, for any $\mathsf{GNN}$ ${\mathcal{M}}\in\mathbb{M}$.
An $\mathsf{IFGC}$ aims to generate a compressed graph $G_{c}$ with a smaller size, such that an inference query that requests output ${\mathcal{M}}(G,V_{T})$ for any $V_{T}\subseteq V$ can be computed by a faster inference process of ${\mathcal{M}}$ over $G_{c}$ only, even with a query-time overhead incurred by ${\mathcal{P}}$.
A Sufficient Condition. We next introduce a sufficient condition for the existence of $\mathsf{IFGC}$. To this end, we start with a notion of inference-equivalent relation.
Inference equivalence. Given a class of $\mathsf{GNN}$ $\mathbb{M}^{L}$ and a graph $G$, a pair of nodes $(v,v^{\prime})$ in $G$ are inference equivalent w.r.t. $\mathbb{M}^{L}$, denoted as $v\sim^{L}_{M}v^{\prime}$, if for any $M\in\mathbb{M}^{L}$, $X_{v}^{k}$ = $X^{k}_{v^{\prime}}$ for any $k\in[1,L]$.
One can readily infer that for any two nodes $v\sim^{L}_{M}v^{\prime}$, $M(v,G)$ = $M(v^{\prime},G)$. That is, inference equivalence of nodes ensure that they are all “indistinguishable” for the inference of any $\mathsf{GNN}$ ${\mathcal{M}}\in\mathbb{M}^{L}$.
Denote the binary relation $(v,v^{\prime})$ induced by inference equivalence as $R^{L}_{M}$, i.e., $(v,v^{\prime})\in R^{L}_{M}$ if and only if $v\sim^{L}_{M}$. We say $R^{L}_{M}$ is nontrivial if there is at least one pair $(v,v^{\prime})\in R^{L}_{M}$, where $v\neq v^{\prime}$. We can readily verify the following result.
Lemma 1.
Given $\mathbb{M}$ and $G$, the binary relation $R^{L}_{M}$ is an equivalence relation, i.e., it is reflexive, symmetric, and transitive.
The equivalent class of $v$ under an equivalence relation $R^{L}_{M}$, denoted as $[v]$, refers to the set $\{v^{\prime}|(v,v^{\prime})\in R^{L}_{M}\}$. The equivalent classes induced by the inference equivalence relation $R^{L}_{M}$ forms a node partition $V_{R}$ of $V$. The quotient graph induced by $R^{L}_{M}$ is a graph $G_{R}$ with nodes $V_{R}$ and edges $E_{R}$, where each node in $V_{R}$ is a distinct equivalent class induced by $R^{L}_{M}$, and there is an edge $([v],[v^{\prime}])\in E_{R}$ if and only if there exists a node $v\in[v]$ and $v^{\prime}\in[v^{\prime}]$, such that $(v,v^{\prime})\in E$.
Lemma 2.
Given a class of $\mathsf{GNNs}$ $\mathbb{M}^{L}$ and a graph $G$, a graph compression scheme $({\mathcal{C}},{\mathcal{P}})$ is an $\mathsf{IFGC}$ w.r.t. $\mathbb{M}^{L}$ and $G$, if for any ${\mathcal{M}}\in\mathbb{M}^{L}$, (1) ${\mathcal{C}}(G)$ computes a quotient graph $G_{c}$ induced by a nontrivial inference equivalent relation $R^{L}_{M}$ w.r.t. $\mathbb{M}^{L}$ and $G$, and (2) ${\mathcal{P}}$ is a function that restores $X^{k}_{v}$ with $X^{k}_{[v]}$ by a scaling factor derived from auxiliary information of $v$, for each layer $k\in[1,L]$.
Proof sketch:  Let $R^{L}_{M}$ be a non-empty inference equivalent relation w.r.t. $\mathbb{M}^{L}$ and $G$, and $G_{c}$ be the quotient graph induced by $\mathbb{M}^{L}$. (1) Given Lemma 1, $R^{L}_{M}$ is a nontrivial equivalence relation. Hence there exists at least one equivalent class $[v]$ with size larger than one, i.e., $|G_{c}|\textless|G|$. As function ${\mathcal{P}}$ does not introduce new node or edge to $G_{c}$, we have $|{\mathcal{C}}(G))|=|G_{c}|\textless|{\mathcal{P}}(G_{c})|\textless|G|$. (2) To see ${\mathcal{M}}(G)$ = ${\mathcal{M}}({\mathcal{P}}({\mathcal{C}}(G)))$, i.e., $G_{c}$ preserves inference result, it suffices to show that for every node $v\in G$, ${\mathcal{M}}(G,\{v\})$ = $M({\mathcal{P}}({\mathcal{C}}(G)),\{[v]\})$. This is ensured by (a) the fixed deterministic inference process that applies the same node update function $M_{v}$, and (b) ${\mathcal{P}}$ restores $X^{k}_{v}$ with only $X^{k}_{[v]}$ and a scaling factor, for any layer $k\in[1,L]$. We list examples of ${\mathcal{P}}$ and scaling factors for mainstream $\mathsf{GNNs}$ in Table 4. Hence $({\mathcal{C}},{\mathcal{P}})$ is an $\mathsf{IFGC}$. $\Box$
We next introduce practical $\mathsf{IFGC}$ for representative $\mathsf{GNN}$ classes, with efficient compression (implementing ${\mathcal{C}}$) and inference (involving ${\mathcal{P}}$) algorithms. We summarize notations in Table 2.
4. Structural-preserving Compression
We introduce a first $\mathsf{IFGC}$ for $\mathsf{GNN}$ inference. We specify $R^{L}_{M}$ as an extended version of structural equivalence. The latter has origins in role equivalence in social science (Lorrain and White, 1971), and simulation equivalence of Kripke structures in model checking (Aziz et al., 1994; Dovier et al., 2001). By enforcing equivalence on embeddings and neighborhood connectivity, it ensures an $\mathsf{IFGC}$ to accelerate $\mathsf{GNN}$ inference without decompression.
4.1. Compression Scheme
Structural equivalence. Given a graph $G$=$(X,A)$, a structural equivalence relation, denoted as $R^{S}$, is a non-empty binary relation such that for any node pair $(v,v^{\prime})$ in $G$, $(v,v^{\prime})\in R^{S}$, if and only if:
  • $\circ$
    $X^{0}_{v}$ = $X^{0}_{v^{\prime}}$, i.e., $v$ and $v^{\prime}$ have the same input features;
  • $\circ$
    for any neighbor $u$ of $v$ ($u\in N(v)$), there exists a neighbor $u^{\prime}$ of $v^{\prime}$ ($u^{\prime}$ in $N(v^{\prime})$), such that $(u,u^{\prime})\in R^{S}$; and
  • $\circ$
    for any neighbor $u^{\prime\prime}$ of $v^{\prime}$ in $N(v^{\prime})$, there exists a neighbor $u^{\prime\prime\prime}$ of $v$ in $N(v)$, such that $(u^{\prime\prime},u^{\prime\prime\prime})\in R^{S}$.
Structural-preserving Compression. Given a $\mathsf{GNN}$ class $\mathbb{M}^{L}$ and graph $G$, a structural-preserving compression, denoted as $\mathsf{SPGC}$ w.r.t. $\mathbb{M}^{L}$, is a pair $({\mathcal{C}},{\mathcal{P}})$ where (1) ${\mathcal{C}}$ computes $G_{c}$ as the quotient graph of $R^{S}$, where $R^{S}$ is the non-empty, maximum structural equivalence relation in $G$, and (2) ${\mathcal{P}}$ is a function that restores node embeddings with matching scaling factors for $\mathbb{M}^{L}$.
Example 1.
Consider the graphs $G_{2}$ and $G_{3}$ in Fig. 2, and their compressed counterpart obtained by $\mathsf{SPGC}$, $G_{c_{2}}$ and $G_{c_{3}}$, respectively.
(1) $G_{2}$ has $10$ nodes and $10$ edges. The nodes having the same labels ${}^{\prime}a^{\prime}$, ${}^{\prime}c^{\prime}$, ${}^{\prime}d^{\prime}$ also have the same input features, respectively. For example, $X^{0}_{a_{1}}$ = $X^{0}_{a_{2}}$, and $X^{0}_{d_{1}}$ = $X^{0}_{d_{2}}$ = $X^{0}_{d_{3}}$. For nodes labeled with ${}^{\prime}b^{\prime}$, $X^{0}_{b_{1}}=X^{0}_{b_{3}}\neq X^{0}_{b_{2}}$. One can verify that $R^{S}$ = $\{(b_{1},b_{3}),(c_{1},c_{2}),(d_{1},d_{3})\}$. A compressed graph $G_{c_{2}}$ is illustrated with $7$ nodes and $7$ edges.
Observe that despite $d_{1}$, $d_{2}$ and $d_{3}$ have the same input features, $d_{2}\not\sim^{L}_{M}d_{1}$, and $d_{2}\not\sim^{L}_{M}d_{3}$ for $\mathsf{GNNs}$ with $L\geq 1$. Indeed, $d_{2}$ has a neighbor $b_{2}$ that has no counterpart in the neighbors of $d_{1}$ or $d_{3}$ that share the same embedding; hence the output embedding of $d_{2}$ may be different from either $d_{1}$ or $d_{3}$, and should be separated from equivalent class $\{d_{1},d_{3}\}$ in $G_{c}$.
(2) $G_{3}$ is a cycle in the form of $\{c_{n},b_{n},a_{n},\ldots,c_{1},b_{1},a_{1}\}$, where $X^{0}_{a_{i}}$ = $X^{0}_{a_{j}}$, $X^{0}_{b_{i}}$ = $X^{0}_{b_{j}}$, and $X^{0}_{c_{i}}$ = $X^{0}_{c_{j}}$, for any $i,j\in[1,n]$. We can verify that $R^{S}=\bigcup_{i,j\in[1,n]}\{(a_{i},a_{j}),(b_{i},b_{j}),(c_{i},c_{j})\}$. A smallest compressed graph $G_{c_{3}}$ is illustrated with only three nodes: $A$ = $\{a_{i}\}$, $B$ = $\{b_{i}\}$, $C$ = $\{c_{i}\}$, $\forall i\in[1,n]$, regardless of how large $n$ is.
The result below tells us that any two nodes that are structural equivalent are “indistinguishable” for $\mathsf{GNN}$ inference process.
Theorem 2.
Given a class of $\mathsf{GNNs}$ $\mathbb{M}^{L}$ and graph $G$, the relation $R^{S}$ over $G$ is an inference equivalence relation w.r.t. $\mathbb{M}^{L}$.
Proof sketch:  First, $R^{S}$ is an equivalence relation. It then suffices to show that for any pair $(v,v^{\prime})\in R^{S}$, $v\sim^{L}_{M}v^{\prime}$. We perform an induction on the number of layers $k$ for $\mathsf{GNNs}$. Consider a “matching” relation $h$ between a pair $(v,v^{\prime})\in R^{S}$, such that $h(v)$ = $v^{\prime}$. At any layer, for any node $v$ and every neighbor $u\in N(v)$, there exists a “match” $h(v)$ and a “match” $h(u)\in N(h(v))$ with the same (intermediate) embedding. This ensures the equivalence of aggregated embedding computed by the node update function at $v$ and $h(v)$, and vice versa. Hence for any pair $(v,v^{\prime})\in R^{S}$, $v\sim^{L}_{M}v^{\prime}$. $R^{S}$ is thus an inference-friendly relation by definition. $\Box$
Following Lemma 2 and Theorem 2, $\mathsf{SPGC}$ is an $\mathsf{IFGC}$.
Example 3.
We also compare $\mathsf{SPGC}$ with several other possible graph compression scheme. We illustrate three more compressed graphs, $G^{1}_{c2}$, $G^{2}_{c2}$, $G^{3}_{c2}$ of $G_{2}$, following Exact Compression (Bollen et al., 2023), Bisimulation (Dovier et al., 2001), and Automorphism (Xu et al., 2021), respectively.
(1) Exact Compression (Bollen et al., 2023) applies an iterative color refinement process [^2] The original method is used to simplify $\mathsf{GNNs}$ learning problems (Bollen et al., 2023); we make a comparison by applying color refinement for graph compression alone. starting with groups that contains nodes agreeing on embeddings and color encoding (labels). It then iteratively split the groups, where each node is updated by concatenating its color encoding with those from their neighbors, and refine groups. It finally derives $G^{1}_{c_{2}}$ with $8$ nodes and $9$ edges after two rounds, where only node pairs $d_{1},d_{3}$ and $a_{1},a_{2}$ share the same node representation. The concatenation is more sensitive to the impact of e.g., degrees, preventing more possible merge, hence less can be compressed.
(2) Bisimulation (Dovier et al., 2001) ignores embedding equivalence and groups nodes only with topology-level equivalence. In this case, $b_{2}$ can be merged with $b_{1}$ and $b_{3}$ due to bisimulation in connectivity. This leads to smaller compressed graph $G^{2}_{c_{2}}$ with $4$ nodes and $3$ edges compared with $G_{c_{2}}$, yet at the cost of inaccurate inference at e.g., nodes $c_{1}$, $c_{2}$, $b_{2}$, and $d_{2}$ due to “overly” compressed structure.
(3) Automorphism (Xu et al., 2021) partitions nodes into same automorphism equivalence sets, which poses strong topological equivalence on graph isomorphism in their neighbors. By enumerating the automorphism groups of $G_{2}$ and considering the embedding similarities, only $a_{1}$ and $a_{2}$ can be merged as shown in $G^{3}_{c2}$ while other node pairs like $c_{1},d_{1}$ or $b_{1},b_{3}$ cannot be merged due to different embeddings or different connection patterns of the neighbors. This can be an overkill for reducing unnecessary inference computation. In addition, computing automorphism remains $\mathsf{NP}$-hard, which indicates more expensive compression cost; while $\mathsf{SPGC}$ computes maximum structural equivalence in $\mathsf{PTIME}$ (see Section 4.3).
Figure 2. Compressing graphs with $\mathsf{SPGC}$ and a comparison with exact compression (Bollen et al., 2023) , bisimulation (Dovier et al., 2001), and automorphism (adapted from  (Xu et al., 2021)). Exact compression uses node features for initial coloring and $d=2$ for color refinement.
4.2. Properties and Guarantees
We next justify $\mathsf{SPGC}$ by showing a minimality and uniqueness property. We show that an $\mathsf{SPGC}$ can generate a smallest $G_{c}$, which is “unique” up to graph isomorphism. That is, if there is another smallest compressed graph $G_{c}^{\prime}$ generated by an $\mathsf{SPGC}$, then $G^{*}_{c}$ and $G_{c}^{\prime}$ are isomorphic.
Lemma 4.
Given a $\mathsf{GNN}$ class $\mathbb{M}^{L}$ and $G$, there is an $\mathsf{SPGC}$ that computes a smallest $G^{*}_{c}$ which is unique up to graph isomorphism.
Proof sketch:  We show the minimality property by a constructive proof as follows: (1) given $\mathbb{M}^{L}$ and $G$, there exists a unique, largest inference equivalence relation $R^{*S}$; (2) we construct an $\mathsf{SPGC}$ that computes $G^{*}_{c}$ as the quotient graph of $R^{*S}$. The uniqueness of $G^{*}_{c}$ can be shown by a contradiction: if there exists another smallest $G_{c}^{\prime}$ that is not graph isomorphic to $G^{*}_{c}$, then either $G_{c}^{\prime}$ is not smallest in sizes, or $R^{*S}$ is not the (largest) inference equivalence relation, i.e., there is a pair $(v,v^{\prime})$, such that either $v\sim^{R}_{M}v^{\prime}$ but are not in $[v]$, or $v\not\sim^{R}_{M}v^{\prime}$, but are included in $[v]$. Either leads to contradiction. $\Box$
We next justify $\mathsf{SPGC}$ by showing that it properly preserves the discriminative set of $\mathsf{GNNs}$, which has been used as one way to characterize the expressiveness power of $\mathsf{GNNs}$ as queries (Barceló et al., 2020; Geerts, 2023).
[…]
Given a set of graphs ${\mathcal{G}}$, denote the set of corresponding compressed graphs generated by $\mathsf{SPGC}$ as ${\mathcal{G}}_{c}$, i.e., ${\mathcal{G}}_{c}$ = $\{G_{c}|G_{c}={\mathcal{C}}(G)$; $G\in{\mathcal{G}}\}$. We have the following result.
Lemma 5.
Given $\mathbb{M}^{L}$ and a set of graphs ${\mathcal{G}}$, an $\mathsf{SPGC}$ can compute a compressed set ${\mathcal{G}}_{c}$, such that for every $\mathsf{GNN}$ ${\mathcal{M}}\in\mathbb{M}^{L}$, and any pair $(G,G^{\prime})\in{\mathcal{G}}_{\mathcal{M}}$, there exists a pair $(G_{c},G^{\prime}_{c})\in{\mathcal{G}}_{c_{\mathcal{M}}}$.
This result tells us that $\mathsf{SPGC}$ “preserves” the discriminativeness of $\mathsf{GNNs}$. Moreover, it suggests a practical compression scheme for large-scale graph classification. One can apply $\mathsf{SPGC}$ to compress ${\mathcal{G}}$ to a smaller counterpart ${\mathcal{G}}_{c}$. As the discriminativeness set is preserved over ${\mathcal{G}}_{c}$ for every $\mathsf{GNN}$ $M\in{\mathcal{M}}^{L}$, $\mathsf{SPGC}$ reduces the overall classification overhead, via a post-processing ${\mathcal{P}}$ that readily groups ${\mathcal{G}}$ by corresponding label groups over ${\mathcal{G}}_{c}$.
Due to limited space, we present the detailed proofs in (ful, 2025).
Algorithm 1 : $\mathsf{SPGC}$
0:  Graph $G$, node feature matrix $X$, a class of $\mathsf{GNNs}$ $\mathbb{M}^{L}$ with node update function $M_{v}$;
0:  A compressed graph $G_{c}$ with memoization structure ${\mathcal{T}}$;
1:  set $R^{S}$ := $\emptyset$; set $EC$ := $\{V\}$; set ${\mathcal{T}}$ := $\emptyset$; graph $G_{c}$ := $\emptyset$;
2:  $R^{S}$ := ${\mathsf{DPP}}(G)$;
3:  $R^{S}$ := $R^{S}\setminus\{(v,v^{\prime})$ |||| $X^{0}_{v}\neq X_{v}^{\prime 0}\}$;
4:  $EC$ := $V/R^{S}$; /* induce partition $EC$ from refined $R^{S}$*/
5:  $(G_{c},{\mathcal{T}})$ := $\mathsf{CompressG}$ (${\mathcal{T}}$, $G_{c}$, $EC$, $G$, $M$);
6:  return  $G_{c}$ and ${\mathcal{T}}$;
Figure 3. Algorithm $\mathsf{SPGC}$
4.3. Compression Algorithm
We next present a compression algorithm (function ${\mathcal{C}}$) in $\mathsf{SPGC}$.
General idea. The algorithm, simply denoted as $\mathsf{SPGC}$, follows Lemma 4 to construct the smallest $G^{*}_{c}$ induced by the maximum structure equivalence relation $R^{*}_{S}$. To ensure efficient inferences that only refer to $G_{c}$ without decompression, it (1) uses a memoization structure ${\mathcal{T}}$ to cache the neighborhood statistics specified by node update function $M_{v}$, and (2) rewrites $M_{v}$ to an equivalent counterpart $M_{[v]}$ (see Table 4 for examples), such that the inference can directly process on each $[v]$ in $G_{c}$, and “looks up” ${\mathcal{T}}$ at runtime, to obtain the embeddings for all the nodes in $[v]$, in a single batch.
Compression Algorithm. The $\mathsf{SPGC}$ algorithm, as illustrated in Fig. 3, takes as input a featurized input $G$ = $(X,A)$ and a $\mathsf{GNN}$ class $\mathbb{M}^{L}$ with node update function $M_{v}$. (1) It first extends Dovier-Piazza-Policriti (${\mathsf{DPP}}$) algorithm (Dovier et al., 2001) to compute the maximum structural equivalence relation $R_{S}^{*}$, by enforcing embedding equivalence as an additional equivalence constraint (lines 2-4). This induces a set of equivalence classes $EC$ (a node partition). It then invokes a procedure $\mathsf{CompressG}$ to construct $G_{c}$ as the quotient graph of $R_{S}^{*}$, as well as the memoization structure ${\mathcal{T}}$ (line 5).
Algorithm 2 Procedure $\mathsf{CompressG}$(${\mathcal{T}}$, $G_{c}$, $EC$, $G$, $M$)
1:  for $[v]\in EC$ do
2:     $V_{c}$ = $V_{c}\cup\{[v]\}$;
3:     initialize $[v]_{T}$; /* with row pointers as $v\in[v]$*/
4:  for edge $(u,v)\in E$ do
5:     $E_{c}$ = $E_{c}\cup\{([u],[v])\}$ |||| $u\in[u],v\in[v]$;
6:     if $M.\phi$ is topology sensitive then
7:        $[v]_{T}(v,[u])\mathrel{+}=\frac{1}{\sqrt{deg(u)}}$;
8:     else if $M.\phi$ is weight sensitive then
9:        $[v]_{T}(v,[u])\mathrel{+}=\alpha_{v,u}$;
10:     else
11:        $[v]_{T}(v,[u])\mathrel{+}=1$;
12:  ${\mathcal{T}}=\bigcup_{[v]\in V_{c}}[v]_{T}$;
13:  return  $G_{c}$ and ${\mathcal{T}}$;
Figure 4. Procedure $\mathsf{CompressG}$
Procedure $\mathsf{CompressG}$. $\mathsf{CompressG}$ is a light-weighted compression approach enabling an easy-to-implement and feasible way to derive compressed graph $G_{c}$ from large-scale $G$ while computing the memoization structure ${\mathcal{T}}$. Given the induced equivalence classes $EC$ and an encoding of the node update function $M_{v}$, procedure $\mathsf{CompressG}$ (illustrated in Fig. 4), generates the compressed graph $G_{c}$ and memoization structure ${\mathcal{T}}$. For each equivalent class $[v]$ in $EC$, $\mathsf{CompressG}$ initializes a node in $G_{c}$, (lines 1-3). For each edge $(u,v)\in E$, $\mathsf{CompressG}$ adds an edge $([u],[v])$ (lines 4-5).
Compression time memoization. $\mathsf{CompressG}$ dynamically maintains a memoization structure ${\mathcal{T}}$ that is shared by all $\mathsf{GNNs}$ in ${\mathcal{M}}^{L}$, to cache useful auxiliary neighborhood information used by the node update function for efficient inference (see Section 4.4). For each node $[v]\in V_{c}$, it assigns $[v]_{\mathcal{T}}$, a compact table, such that for every $v\in[v]$, and every neighbor $[u]\in N([v])$, an entry $[v]_{{\mathcal{T}}}(v,[u])$ records an aggregation of auxiliary neighborhood information (e.g., sum of node degree, edge weights) of $N(v)\subseteq N([v])$.
When processing an edge $(u,v)\in E$, it follows a case analysis of $\mathbb{M}^{L}$ with node update function $M_{v}$. For example, (1) “topology sensitive” means the degrees of neighbors of $v$ is required, as seen in $\mathsf{GCNs}$; (2) “weight sensitive” means additional edge weights, such as edge attentions in $\mathsf{GATs}$. For $\mathsf{GATs}$, the edge weights from pre-trained $M$ are optional such that if given the information of all edge weights, the inference accuracy can be preserved based on the formula shown in Table. 4.4. Such information can be readily obtained by tagging the input $\mathsf{GNN}$ class $\mathbb{M}^{L}$ or encoded as rules. It then updates the entry $[v]_{{\mathcal{T}}}(v,[u])$ accordingly (lines 6-12).
Example 6.
Consider a $\mathsf{GNNs}$ class $\mathsf{GIN}$, and graph $G_{4}$ shown in Fig. 5. (1) $\mathsf{SPGC}$ invokes the ${\mathsf{DPP}}$ algorithm to compute $R^{S}$. It next refines $R^{S}$ based on feature embeddings and returns the induced partition $EC=\{[a],[b],[c],[d]\}$, where nodes with same labels are merged, e.g., $[a]=\{a_{1},a_{2}\}$. (2) We illustrate how $\mathsf{CompressG}$ dynamically updates the memoization structure ${\mathcal{T}}$ by considering the processing of two edges $(b_{1},a_{1})$ and $(b_{2},a_{1})$. For $[a]\in EC$, it first initializes $[a]_{\mathcal{T}}$ as an empty table. It next iterates over edges in $E$. As the node update function of $\mathsf{GIN}$ does not require exact degree or additional edge weight (not topology or weight sensitive), the entry $[a]_{T}(a_{1},[b])$ is updated to 1, to “memoize” that there is one neighbor of $a_{1}$ in $N([b])$ that will contribute to a “unit” value to the embedding computation of $a_{1}$, via edge $(b_{1},a_{1})$. Similarly, when it reaches edge $(b_{2},a_{1})$, $[a]_{T}(a_{1},[b])$ is updated to 2. Following this processes, all entries in ${\mathcal{T}}$ will be updated to memoize neighbors’ information while compressing the graph.
Figure 5. Run-time generation of Memoization structure ${\mathcal{T}}$.
Correctness and cost. $\mathsf{SPGC}$ correctly computes $G^{*}_{c}$ as ensured by (1) the correctness of DPP algorithm and (2) the follow-up refinement by enforcing embedding equivalence. For time cost, it takes $\mathsf{SPGC}$ $O(|V|+|E|)$ time to initialize $R^{S}$ with DPP algorithm. The refinement of $R^{S}$ and $EC$ takes $O(|V|)$ time (lines 3-4). Procedure $\mathsf{CompressG}$ processes each equivalent class in $EC$ ($|EC|\leq|V|$) and each edge in $G$ once, hence in $O(|V|+|E|)$ time to construct $G^{*}_{c}$ and update ${\mathcal{T}}$. The total cost is thus in $O(|V|+|E|)$ time.
$\mathsf{GNNs}$ Node Update Function $M_{v}$ equivalent rewriting $M_{[v]}$; scaling factors are marked in red notes
Vanilla (Scarselli et al., 2008) $X_{v}^{k}=\sigma(\Theta\cdot\text{AGG}(X_{u}^{k-1},\forall u\in\mathcal{N}(v)))$ $X_{v}^{k}=\sigma(\Theta\cdot\text{AGG}(\mathbf{{\color[rgb]{1,0,0}[v]_{T}(v,[u])}}X_{[u]}^{k-1},\forall[u]\in\mathcal{N}([v])))$ AGG: $\sum$ or ${\mathsf{AVG}}$; for $AVG$, need to multiple by $RF_{v}$
$\mathsf{GCN}$ (Kipf and Welling, 2016) $X_{v}^{k}=\sigma(\Theta^{k}(\sum_{u\in\mathcal{N}(v)}\frac{1}{\sqrt{deg_{u}deg_{v}}}x_{u}^{k-1}))$ $X_{v}^{k}=\sigma(\Theta^{k}(\sum_{[u]\in\mathcal{N}([v])}\mathbf{\color[rgb]{1,0,0}\frac{1}{\sqrt{deg_{v}}}[v]_{T}(v,[u])}x_{[u]}^{k-1}))$ $deg_{v}$: degree of node $v$ in $G$, topology sensitive
$\mathsf{GAT}$ (Velickovic et al., 2017) $X_{v}^{k}=\sigma(\sum_{u\in\mathcal{N}(v)}\alpha_{uv}\Theta^{k}X_{u}^{k-1}))$ $X_{v}^{k}=\sigma(\sum_{[u]\in\mathcal{N}([v])}\mathbf{{\color[rgb]{1,0,0}[v]_{T}(v,[u])}}\Theta^{k}X_{[u]}^{k-1}))$ weight sensitive
$\mathsf{GraphSAGE}$ (Hamilton et al., 2017) $X_{v}^{k}=\sigma(\Theta^{k}\cdot(X_{v}^{k-1}||\text{AGG}(X_{u}^{k-1},\forall u\in\mathcal{N}(v))))$ $X_{v}^{k}=\sigma(\Theta^{k}\cdot(X_{[v]}^{k-1}||\text{AGG}(\mathbf{{\color[rgb]{1,0,0}RF_{v}\times[v]_{T}(v,[u])}}X_{[u]}^{k-1}$, $\forall[u]\in\mathcal{N}([v]))))$ ||||| |: concatenation; AGG: ${\mathsf{AVG}}$
$\mathsf{GIN}$ (Xu et al., 2019) $X_{v}^{k}=\sigma(\text{MLP}((1+\gamma)x_{v}^{k-1}+\sum_{u\in\mathcal{N}(v)}x_{u}^{k-1}))$ $X_{v}^{k}=\sigma(\text{MLP}((1+\gamma)x_{[v]}^{k-1}+\sum_{[u]\in\mathcal{N}([v])}\mathbf{{\color[rgb]{1,0,0}[v]_{T}(v,[u])}}x_{[u]}^{k-1}))$
Table 4. Rewriting of node update functions for mainstream $\mathsf{GNN}$ classes (scaling factors highlighted in red).
4.4. Inference Process
Inference algorithm. We outline an algorithm that directly obtains $M(G)$ by referring to $G^{*}_{c}$ only, without decompression. Our strategy rewrites the node update function $M_{v}$ to an equivalent counterpart $M_{[v]}$, that takes as input $[v]$ and the corresponding tuple $[v]_{\mathcal{T}}(v)$ in ${\mathcal{T}}$, to “scale” the embedding computation with the memorized edge weights. The algorithm performs inference directly in $G^{*}_{c}$ with $M_{[v]}$, and simply “scale up” the results at $[v]$ for each node $v\in[v]$ with a scaling factor. The scaling factor can be directly looked up from the table $[v]_{\mathcal{T}}(v)$ (function ${\mathcal{P}}$). Table 4 illustrates the scaling factors for mainstream $\mathsf{GNN}$ classes.
Example 7.
Continuing with Example. 6, an inference at node $a_{1}$ looks up, in constant time, the values from entries $[a]_{T}(a_{1},[b])$ and $[a]_{T}(a_{1},[c])$ which are 2 and 1 separately (as shown in Fig. 5). It next assigns the values as scaling factors (Table. 4) to restore messages and the embedding of node $a_{1}$ as in original $G_{4}$.
Inference cost. As $\mathsf{SPGC}$ requires no decompression on neighborhood structures of nodes, an inference query can be directly applied to $G_{c}$ without incurring additional overhead. The overall inference cost is in $O(L|E_{c}|F+L|V_{c}|F^{2})$. We remark that this result is derived by scaling down a common upper bound of inference costs for mainstream $\mathsf{GNNs}$ in Table 3. For other and more complex $\mathsf{GNNs}$ variants, the inference costs can be derived similarly by scaling down from their counterparts over $G$.
5. Configurable Graph Compression
While $\mathsf{SPGC}$ generates $G_{c}$ that can be directly queried by inference queries without decompression, it enforces node embedding equivalence, which may be an overkill for nodes with similar embeddings and can be processed in a single batch with tolerable difference in query outputs. Users may also want to configure the compression schemes to balance among accuracy and speed up, or to contextualize the compression with inference queries that specifies a set of test nodes $V_{T}\subseteq V$ of interests, such that $M(G,V_{T})$ = $M(G_{c},V_{T})$.
In response, we next introduce two variants of $\mathsf{SPGC}$: $(\alpha,r)$-$\mathsf{SPGC}$ (Section 5.1), and anchored $\mathsf{SPGC}$ (Section 5.2), respectively.
5.1. Compression with Structural and Embedding Similarity
We start with a relation called $(\alpha,r)$-relation, which approximates $R^{S}$ by lifting its equivalence constraints.
$(\alpha,r)$-relation. Given graph $G$, a configuration $({\mathsf{xsim}},\alpha,r)$ is a triple where ${\mathsf{xsim}}(\cdot)$ is a feature similarity function that computes a similarity score of two node embeddings, $\alpha$ is a similarity threshold ($\alpha\in[0,1]$), and $r$ an integer. Let $N_{r}(v)$ be the nodes within $r$-hop neighbors of $v$. A binary relation $R^{(\alpha,r)}\subseteq V\times V$ is an $(\alpha,r)$-relation if for any node pair $(v,v^{\prime})\in R^{(\alpha,r)}$,
  • $\circ$
    ${\mathsf{xsim}}(X_{v}^{0},X_{v^{\prime}}^{0})\geq\alpha$;
  • $\circ$
    for any node $u\in N(v)$, there exists a node $u^{\prime}\in N_{r}(v^{\prime})$), such that $(u,u^{\prime})\in R^{(\alpha,r)}$; and
  • $\circ$
    for any node $u^{\prime\prime}\in N(v^{\prime})$, there exists a node $u^{\prime\prime\prime}\in N_{r}(v)$, such that $(u^{\prime\prime},u^{\prime\prime\prime})\in R^{(\alpha,r)}$.
Figure 6. Compressing $G_{5}$ with $(0.5,2)$-$\mathsf{SPGC}$.
Note that $R^{(1,1)}$ is an $R^{S}$, as $\alpha$ = 1 ensures embedding equivalence, and $r$ = 1 preserves indistinguishable neigbhors for node update functions in $\mathsf{GNN}$ inference. On the other hand, $(\alpha,r)$-relation is no longer an equivalence relation, as it “relaxes” structural equivalence by lifting both embedding equality, and the strict neighborhood-wise equivalence, in trade for better compression ratio. We next clarified the relationship between $r$ and $L$. Based on the definition of $\mathsf{SPGC}$, the embedding similarity between two vertices is dependent on $r$ instead of $L$. In other words, $\mathsf{SPGC}$ is model-agnostic as long as $r\leq L$. In practice, since $L$ is small due to over-smoothing issue of larger $L$, $r$ is usually a small integer like 1 or 2.
Based on the relation $R^{(\alpha,r)}$, we introduce a variant of $\mathsf{SPGC}$.
$(\alpha,r)$-$\mathsf{SPGC}$. Given a graph $G$, and a configuration $\alpha$ and $r$ w.r.t. an embedding similarity measure and a threshold, an $(\alpha,r)$-$\mathsf{SPGC}$ is a graph compression scheme if ${\mathcal{C}}$ computes a graph $G_{c}$ induced by the relation $R^{(\alpha,r)}$. Specifically,
  • $\circ$
    for any node pair $(v,v^{\prime})\in R^{(\alpha,r)}$, $v\in[v]$, $v^{\prime}\in[v]$; and
  • $\circ$
    there is an edge between $([u],[v])$ if $(u,v)\in E$.
Lemma 1.
Given a $\mathsf{GNN}$ class $\mathcal{M}^{L}$ and a graph $G$, an $(\alpha,r)$-$\mathsf{SPGC}$ incurs compression cost in $O(|V||N_{r}|+|E|)$ time ($|N_{r}|$ refers to the largest size of $r$-hop neighbors of a node in $G$), and an inference cost in $O(L|E|F+L|V_{c}|F^{2})$ time.
As a constructive proof, we next introduce algorithms that implements $(\alpha,r)$-$\mathsf{SPGC}$ with the above guarantees.
Compression Algorithm. We describe the compression algorithm $(\alpha,r)$-$\mathsf{SPGC}$. It follows the same principle to compute $(\alpha,r)$-relation and induce a compressed graph. The difference is that rather than inducing equivalence class and quotient graph from $G$, (1) it first induces a graph $G_{r}$ by linking nodes to their $r$-hop neighbors, (2) it then computes an $R^{S}$ relation by invoking $\mathsf{DPP}$ algorithm, and refines it by a re-grouping of nodes determined by similarity function ${\mathsf{xsim}}(\cdot)$ with $\alpha$ as similarity threshold. (3) It generates $[v]$ to include all the pairs $(v,v^{\prime})$ in $R^{(\alpha,r)}$, and accordingly the edges. It updates the memoization structure ${\mathcal{T}}$ following the edges in $G_{r}$, similarly as in $\mathsf{SPGC}$. Here ${\mathcal{T}}$ caches the statistics from the $r$-hop neighbors of each node $v$ in the original graph $G$.
Example 2.
Consider graph $G_{5}$ shown in Fig. 6. A $(0.5,2)$-$\mathsf{SPGC}$ invokes DPP to initialize a $(1,1)$-relation, and refines it to $R^{(0.5,2)}=\{(a_{1},a_{2}),(c_{1},c_{2}),(b_{1},b_{2}),(b_{1},b_{3}),(b_{1},b_{4}),(b_{2},b_{3}),(b_{2},b_{4}),(b_{3},b_{4})\}$. This yields a compressed graph $G^{2}_{c_{5}}$ with only $3$ nodes and $2$ edges.
We also illustrate $G^{1}_{c_{5}}$, a compression graph from ${\mathsf{SPGC}}$ for $G_{5}$ (induced by an $R^{S}$ as a $(1,1)$-relation). Due to strictly enforced embedding equivalence, $b_{1}$ and $b_{2}$ cannot be merged, and similarly for $b_{3}$ and $b_{4}$. This yields $G^{1}_{c_{5}}$ with more nodes and edges.
Compression Cost. It takes $O(|V|\cdot|N_{r}(v)|)$ time to derive $G_{r}$ for $v\in V$. It then takes $O(|V|+|E_{r}|)$ to compute and refine $R^{(\alpha,r)}$, for $G_{r}$ with edge set $E_{r}$. Here $|E_{r}|\leq|V|\cdot|N_{r}|$, where $N_{r}$ refers to the largest $r$-hop neighbor set for a node in $G$. Procedure $\mathsf{CompressG}$ constructs $G_{c}$ in $O(|V|+|E_{r}|)$ time, and generates memoization structure ${\mathcal{T}}$ in $O(|E|)$ time. The total cost is thus in $O(|V||N_{r}|+|E|)$.
As $(\alpha,r)$-$\mathsf{SPGC}$ is specified by $R^{(\alpha,r)}$ that approximates an inference-friendly relation, it is no longer an $\mathsf{IFGC}$, hence a direct inference over the $G_{c}$ from it may not preserve the original output. To mitigate accuracy loss, the inference specifies a procedure ${\mathcal{P}}$ to perform run-time decompression with small overhead.
Inference process with decompression. The inference algorithm directly processes each node $[v]$ in $G_{r}$ as in $\mathsf{SPGC}$. For $(\alpha,r)$-$\mathsf{SPGC}$. the difference is that it ad-hocly invokes a decompression procedure $\mathsf{decompG}$ (the $\mathsf{decompG}$ algorithm and its example illustrated in Figure. 18 are shown in the Appendix (ful, 2025)) to reconstruct the neighbors of $v$ in $G_{r}$, and performs an inference using original $1$-hop neighbors of $v$ to obtain an embedding $X_{v}$ as close as its original counterpart in $G_{r}$. To minimize decompression cost, when compressing the graph, algorithm $(\alpha,r)$-$\mathsf{SPGC}$ (shown in Appendix B) incorporates Re-Pair, a reference encoding method (Larsson and Moffat, 2000; Claude and Navarro, 2007), to derive $\mathsf{AL_{c}}$ and $\mathsf{rules}$ for later fast decompression from a compact encoding structure. Specifically, within each $EC$, procedure $\mathsf{decompG}$ sorts the processing order of nodes by their degrees in $G_{r}$. In other words, procedure $\mathsf{decompG}$ prioritizes the decompression of nodes having the most shared $r$-hop neighbors with others in $G_{r}$, to (1) maximally reduce redundant computation in decompression process, and (2)“maximize” the likelihood for more accurate inference computation. For example, a partially decompressed graph $G_{cd_{5}}$ that resolves $1$-hop neighbors of $b_{2}$ (in trade for more accurate embedding) is illustrated in Fig. 6 (please refer to Example. 8 and Figure. 18 in Appendix for details (ful, 2025)). The decompressed neighbors are kept in $G_{cd}$ until the inference terminates.
As the decompression restores at most $|E|$ edges, the overall inference process takes $O(L|E|F+L|V_{c}|F^{2})$ time, including the decompression overhead. We present the details of decompression algorithm in (ful, 2025). The above analysis completes the proof of Lemma 1.
5.2. Anchored Graph Compression
We next introduce our second variant of $\mathsf{SPGC}$, notably, anchored $\mathsf{SPGC}$, which permits a decompression-free, inference friendly compression, relative to a specific set of nodes of interests.
We present our main result below.
Theorem 3.
Given $\mathcal{M}^{L}$ and $G$ with a set of targeted nodes $V_{A}$, there exists an $\mathsf{IFGC}$ that computes a compressed graph in $O(|G_{L}|)$ time to preserve the inference output for every node in $V_{A}$ at an inference cost in $O(L|E_{c}|F+L|V_{c}|F^{2})$ time. Here $|G_{L}|$ refers to the subgraph of $G$ induced by $L$-hop of $V_{A}$, and $|V_{c}|$ and $|E_{c}|$ are bounded by $|G_{L}|$.
As a constructive proof, we introduce a notion of anchored relation, and construct such an $\mathsf{IFGC}$. Given a graph $G$ with a set of designated targeted nodes $V_{A}$, and a class of $\mathsf{GNNs}$ $\mathcal{M}^{L}$, a graph compression scheme $({\mathcal{C}},{\mathcal{P}})$ is an $\mathsf{IFGC}$ relative to $V_{A}$, if (1) $|{\mathcal{P}}({\mathcal{C}}(G))|<|G|$; and (2) for any $\mathsf{GNN}$ $M\in\mathcal{M}^{L}$, and any $v\in V_{A}$, $M(G,\{v\})$ = $M({\mathcal{P}}({\mathcal{C}}(G),\{v\})$.
Anchored relation. Given graph $G$, an integer $L$, and a designated anchor set $V_{A}\subseteq V$, we define the $L$-hop neighbors of $V_{A}$, denoted as $N_{L}(V_{A})$, as $\bigcup_{v\in V_{A}}N_{L}(v)$, where $N_{L}(v)$ refers to the set of nodes within $L$-hop of $v$ in $G$. An anchored relation $R^{A}_{L}$ w.r.t. $V_{A}$ refers to the structural equivalence relation defined over the subgraph $G_{L}$ of $G$ induced by $N_{L}(V_{A})$.
One may verify that (1) $R^{A}_{L}$ is an equivalence relation over $V$, and (2) $R^{S}$ = $R^{A}_{L}$ if $V_{A}$ = $V$, and $L$ is larger than the diameter of $G$.
Anchored $\mathsf{SPGC}$. Given graph $G$ and an anchor set $V_{A}$ from $G$, an anchored $\mathsf{SPGC}$, denoted as $\mathsf{ASPGC}$, is a graph compression where ${\mathcal{C}}$ computes a compressed graph that is the quotient graph of $R^{A}_{L}$. The computation of compressed graph $G_{c}$ using $\mathsf{ASPGC}$ follows its counterpart in $\mathsf{SPGC}$. The difference is that it induces a subgraph $G_{L}$ of $G$ with the $L$-hop neighbors of all the nodes in $V_{A}$. It then invokes the compression algorithm of $\mathsf{SPGC}$ to derive $R^{A}_{L}$ and applies the compression algorithm of $\mathsf{SPGC}$ on $G_{L}$ to compute the $G_{c}$ and memoization structure ${\mathcal{T}}$. The inference process over $G_{c}$, similarly, follows its $\mathsf{SPGC}$ counterpart over $G_{c}$, which consistently leverages ${\mathcal{T}}$ to efficiently recover the auxiliary information of neighborhoods. Hence, it preserves the inference results of the nodes in $V_{A}$ for $\mathsf{GNNs}$ classes with up to $L$ layers. In practice, one may set $V_{A}$ simply as a set of test nodes $V_{T}$ to adapt $\mathsf{ASPGC}$ for specific inference queries. We present details and an example in (ful, 2025).
Analysis. The correctness of $\mathsf{ASPGC}$ follows from the data locality of $L$-layered $\mathsf{GNN}$ inference when $V_{A}$ is specified, which only involves the subgraph $G^{L}$ of $G$ induced by $L$-hop neighbors of $V_{A}$. $\mathsf{ASPGC}$ next follows $\mathsf{SPGC}$ to correctly compute $G_{c}$ from $G_{L}$. For compression cost, it takes $O(|N^{L}(V_{A})|+|E|)$ time to induce $G_{L}$, and $O(|V_{L}|+|E_{L}|)$ time to construct $G_{c}$ from $G_{L}$. (3) The inference cost is consistently $O(L|E_{c}|F+L|V_{c}|F^{2})$, where both $|E_{c}|$ and $|V_{c}|$ are bounded by $|G_{L}|$.
Given the above analysis, Theorem 3 follows.
6. Parallel Inference
$\mathsf{SPGC}$ and its variants go through the following modules and algorithms to parallelize inference, as illustrated in Fig. 7. (1) Coordinator. Given the configuration tuple $(M^{L},V_{T})$ provided by the user, coordinator conducts the following steps: (1) retrieves the corresponding $G$ and ${\mathcal{T}}$ from the compressor; (2) generates parallel joblets $<{\mathcal{T}}_{i},G_{c_{1}},M^{L}>$ and sends them to processor $P_{i}$ for all $v_{i}\in V_{T}$. (2) Compressor. Given a graph $G$ and a class of $\mathsf{GNNs}$ ${\mathcal{M}}^{L}$, the compressor produces the compressed graph $G_{c}$ and the memoization structure ${\mathcal{T}}$. It performs offline “once-for-all” graph compression by leveraging a built-in library of compression methods, including our structural-preserving compression $\mathsf{SPGC}$ as well as the state-of-the-art graph compression approaches (Liu et al., 2023; Kumar et al., 2023). (3) Parallel Inference. Given the joblet $<{\mathcal{T}}_{i},G_{c_{1}},M^{L}>$, the module computes the inference result $M^{L}(v_{i},G_{c_{i}})$ in parallel.
Coordinator. The coordinator optimizes query workload with three components. (1) Partitioner: Upon receiving an inference query $(M^{L},V_{T})$, it initializes a set of processors $\mathcal{P}$. At each $P_{i}\in\mathcal{P}$, it partitions $V_{T}$, the memoization table ${\mathcal{T}}$ to ${\mathcal{T}}_{i}$, and induces $G_{c_{i}}$ accordingly to $v_{i}\in V_{T}$. (2) Load Balancer: It assigns joblets to processors by distributing query workload evenly among the processors, preventing bottlenecks. This approach minimizes the total time cost of inference on the entire set of $V_{T}$. (3) Job Scheduler: It reschedules joblets $<\tau_{i},G_{c_{i}},M^{L}>$ based on estimated workload and current processor status, and allocate computational resources to dynamically rebalance the parallel inference computation.
Figure 7. Parallel Inference and Explainer with a running example: $M^{L}\in{\mathsf{GCN}}^{2}$; $V_{T}=\{a_{1},b_{1}\}$; ${\mathcal{T}}_{1}$, ${\mathcal{T}}_{2}$: partitioned subsets of ${\mathcal{T}}$ for $a_{1}$ and $b_{1}$; $G_{s_{1}}$, $G_{s_{2}}$: explanation graphs for $a_{1}$ and $b_{1}$.
Pipelining. At each processor $P_{i}$, given a joblet $<{\mathcal{T}}_{i},G_{c_{i}},M^{L}>$, it performs the inference for assigned $v_{i}\in V_{T}$ in parallel to generate output $M^{L}(v_{i},G_{c_{i}})$. The inference joblets follow a pipelined parallelism among all processes, to ensure that users receive inference result incrementally, rather than waiting or all inference queries to be evaluated.
Parallel Time Cost. The parallel cost per inference query is $O(\frac{L|G_{c}|F^{2}}{n})$ where $L$ is the number of layers and $F$ is number of features per node in $G_{c}$, and $n$ the number of processors.
7. Experimental Study
Using both real-world graph datasets and large synthetic graphs, we conducted four sets of experiments, to understand (1) effectiveness of our compression methods, in terms of compression ratio, and the trade-off between inference cost and accuracy loss; (2) their efficiency, in terms of the compression cost and inference cost, (3) impact of critical factors, and (4) an ablation study to evaluate the effectiveness of memoization, and decompression overhead.
7.1. Experimental Settings
Datasets. We employ four real-world graph benchmark datasets (summarized in Table. 5): (1) Cora (McCallum et al., 2000), a citation network where nodes represent documents, and edges are citations among the documents. Node features are described by a 0/1-valued word vector indicating the absence/presence of the corresponding word from the dictionary; (2) Arxiv (Hu et al., 2020), a citation network with nodes representing arXiv papers and edges denoting one paper cites another one. The features of each node includes a 128-dimensional feature vector obtained by averaging the embeddings of words in its title and abstract; (3) Yelp (Zeng et al., 2019) is prepared from the raw json data of businesses, comprising a dense network of user-business interactions. Nodes represent users and edges are generated if two users are friends. Node features are the word embeddings derived by reviews of products; and (4) Products (Hu et al., 2020), a product co-purchase network, where nodes represent products sold in Amazon and edges denote products purchased together. Node features are generated by extracting embeddings from the product descriptions followed by a Principal Component Analysis reducing the dimension to 100. The size of test nodes $|V_{T}|=0.05*|V|$ across all datasets. Note that $|V_{A}|=|V_{T}|$ since by default $V_{A}=V_{T}$ in $\mathsf{ASPGC}$.
Besides real-world benchmark datasets, we also generated a large synthetic dataset WS-MAG, by extending a core of real MAG240M network (Hu et al., 2021) (a citation network) with a small world generator  (Watts and Strogatz, 1998).
Dataset |V| |E| # node types # attributes
Cora 2,708 5,429 7 1,433
Arxiv 169K 1.2M 40 128
Yelp 717K 7.9M 100 300
WS-MAG 2M 8M 153 768
Products 2.4M 61.9M 47 100
Table 5. Summary of Datasets.
Graph Neural Networks. We have pre-trained three classes of representative $\mathsf{GNNs}$: ${\mathsf{GCNs}}^{3}$ (Kipf and Welling, 2016), ${\mathsf{GATs}}^{3}$ (Velickovic et al., 2017), and ${\mathsf{GraphSAGE}}^{3}$ (Hamilton et al., 2017), for each dataset. For a fair comparison, (1) for all the datasets, we consider node classification, and (2) all compression methods are applied for the same set of $\mathsf{GNNs}$
Compression Methods. We compare $\mathsf{SPGC}$ and its variants, $(\alpha,r)$-$\mathsf{SPGC}$ and $\mathsf{ASPGC}$, with two state-of-the-art compression methods: (1) $\mathsf{DSpar}$ (Liu et al., 2023), a graph sparsification method that performs edge down-sampling to preserve graph spectral information, and (2) $\mathsf{FGC}$ (Kumar et al., 2023), a latest learning-based graph coarsening approach that learns a coarsened graph matrix and feature matrix to preserve desired graph properties, such as homophily. We are aware of other learning-based approaches, yet they are model-specific and require the learned model parameters. Our work is orthogonal to these methods, and is not directly comparable. When conducing the experiments using $\mathsf{DSpar}$ and $\mathsf{FGC}$, we set the longest waiting limit = 5 hours which means if the compression graph cannot be generated after 5 hours. We simply exclude them since the compression cost is already comparable to training a new $\mathsf{GNNs}$ model.
Evaluation Metrics. Given graph $G$, a class of ${\mathsf{GNNs}}$ $\mathcal{{\mathcal{M}}}^{L}$, a graph compression scheme $({\mathcal{C}},{\mathcal{P}})$ that computes a compressed graph $G_{c}$, and its matching inference process over $G_{c}$, we use the following metrics. (1) For efficiency, we evaluate (a) the time cost of compression, and (b) the speed-up of inference, which is defined by $\frac{T_{MG}}{T_{MC}}$, where $T_{MG}$ refers to the inference time cost over $G$, and $T_{MC}$ represents its counterpart over $G_{c}$. (2) For effectiveness, we report (a) a normalized compression ratio, which is defined as ${\mathsf{ncr}}$ = $1-\frac{|G_{c}|}{|G|}$; Intuitively, it quantifies the fraction of $G$ that is “reduced”: the larger, the better; and (b) the model performance quantified by accuracy and $F1$-score over $G_{c}$. In particular for $\mathsf{Yelp}$, over which the benchmark task is a multi-class node classification, we report micro $F1$-score. We report the average performance of $200$ inference tests, for each $\mathsf{GNN}$ model over each dataset.
Environment. $\mathsf{SPGC}$ and its variants are developed in Python with PyTorch Geometric (Fey and Lenssen, 2019) and BisPy (Andreuzzi, 2021) libraries. All tests are conducted on 4 Intel(R) Xeon(R) Silver 4216 CPU @ 2.10GHz, 128 GB Memeory, 16 cores, and 1 32GB NVIDIA V100 GPU. Our source code, datasets, and a full version of the paper are made available [^3] https://github.com/Yangxin666/SPGC.
7.2. ExperimentalResults
Exp-1: Effectiveness: Accuracy vs. Speed-up. Fig. 8 compares $(0.5,1)$-$\mathsf{SPGC}$ with $\mathsf{DSpar}$ and $\mathsf{FGC}$, in terms of inference speed-up in left figure and inference accuracy/F1-score in the right figure, over all four real datasets. Note that $\mathsf{FGC}$ can only generate results for Cora since its compression cannot be completed within 5-hours for all other datasets. Here a test annotated as “compression method-$\mathsf{GNN}$” refers to the setting that a $\mathsf{GNN}$ inference is applied on a compressed graph generated by the method. (1) $(0.5,1)$-$\mathsf{SPGC}$ outperforms $\mathsf{DSpar}$ and $\mathsf{FGC}$ across all four datasets for all the $\mathsf{GNN}$ classes on inference speed-up. It can improve the inference efficiency better over larger graphs. For example, for Arxiv, $(0.5,1)$-$\mathsf{SPGC}$ achieves a speed-up of 3.4 for inference with $\mathsf{GraphSAGE}$, while $\mathsf{DSpar}$ achieves a speed-up to 1.5. (2) Consistently, we found that $\mathsf{SPGC}$ achieves higher ${\mathsf{ncr}}$ than $\mathsf{DSpar}$ and $\mathsf{FGC}$. For example, for Cora, $\mathsf{SPGC}$ achieves ${\mathsf{ncr}}$ up to 74.5% while $\mathsf{DSpar}$ and $\mathsf{FGC}$ achieves 46.2% and 30.7% respectively. (3) $\mathsf{SPGC}$ outperforms $\mathsf{DSpar}$ and $\mathsf{FGC}$ in terms of its ability to preserve inference results for $\mathsf{GNNs}$. We observe that i) $(0.5,1)$-$\mathsf{SPGC}$ achieves inference accuracy comparable to direct inference on the original graph $G$ across all four datasets (as illustrated in Fig. 8); and ii) $(0.5,1)$-$\mathsf{SPGC}$ retains highest inference accuracy/F1-scores across all datasets and models. For example, for Cora, $(0.5,1)$-$\mathsf{SPGC}$ achieves 0.71 accuracy which outperforms 0.68 and 0.67 achieved by $\mathsf{DSpar}$ and $\mathsf{FGC}$.
Figure 8. Comparison of $(0.5,1)$-$\mathsf{SPGC}$ with the Baselines in Inference Speed-up and Inference Accuracy/F1-score. (the inference accuracy of $\mathsf{ASPGC}$ marked in colored lines, e.g., the red lines show inference accuracy of $\mathsf{ASPGC}$ using $\mathsf{GCN}$; the inference accuracy on original graph $G$ marked in corresponding colors for difference $\mathsf{GNNs}$, e.g., the blue 0.73 indicates the inference accuracy of GAT on original $G$ of Cora).
Exp-2: Effectiveness: Impact of Factors. We first investigate the impact of number of layers $L$, which evaluates whether the quality of $\mathsf{SPGC}$-based compression is affected by the complexity of $\mathsf{GNNs}$ classes. Then we evaluate the performance of configurable compression $(\alpha,r)$-$\mathsf{SPGC}$, in terms of the impact of $\alpha$ and $r$.
Varying Number of Layers. We varied the number of layers of $\mathsf{GNNs}$ from $2$ to $4$ over Arxiv and report its impact on inference speed-up (resp. accuracy) in Fig. 9(a) (resp. 9(b)). (1) $(0.5,1)$-$\mathsf{SPGC}$ consistently outperforms all the baselines in both inference speed-up and accuracy, due to that it preserves the inference results with small compressed graphs. (2) In general, the speed-up of inference achieved by $(0.5,1)$-$\mathsf{SPGC}$ is not sensitive to the number of layers. This verifies our theoretical analysis that it preserves inference results with unique smallest compressed graphs, which is independent of model complexity. (3) While the accuracy of $\mathsf{GNNs}$ drops as the number of layers become larger, in all cases, $(0.5,1)$-$\mathsf{SPGC}$ preserves the accuracy with smallest “gap” compared with other methods, for the same class of $\mathsf{GNNs}$.
(a) Num. Layers v.s. Speed-up
(b) Num. Layers v.s. Accuracy
Figure 9. Varying Num. Layers in GNNs (Arxiv).
Varying $\alpha$. Fixing $r=1$, we varied $\alpha$ from $0.2$ to $1$, and report the results in Figs. 10(a) to  10(d). It tells us the followings.
(1) As $\alpha$ is increased from 0.2 to 1, ${\mathsf{ncr}}$ drops as illustrated in Fig. 10(a). As expected, larger $\alpha$ makes it harder for $(\alpha,r)$-$\mathsf{SPGC}$ to merge nodes that are less close in their representations, leaving more nodes in compressed graphs, hence worsening compression ratio. (2) Consistently for larger $\alpha$ and for all the three $\mathsf{GNNs}$ classes, it is harder for $(\alpha,r)$-$\mathsf{SPGC}$ to improve their inference efficiency due to larger compressed structure $G_{c}$ as shown in Fig. 10(b). We observe consistent results for Yelp and Products (not shown; see (ful, 2025)).
(3) As $\alpha$ increases, the F1-score (resp. accuracy) of the inference results over Yelp (resp. Products) remains insensitive, as shown in Figs.10(c) (resp. 10(d)). Our observation over Arxiv remains consistent, and we omitted it due to limited space. This indicates that $(\alpha,r)$-$\mathsf{SPGC}$ does not lose much on the quality of the inference while significantly improved inference efficiency.
(a) $\alpha$ v.s. ${\mathsf{ncr}}$
(b) $\alpha$ v.s. Speed-up (Arxiv)
(c) $\alpha$ v.s. F1-score (Yelp)
(d) $\alpha$ v.s. Accuracy (Products)
(e) $r$ v.s. Speed-up
(f) $r$ v.s. Accuracy
Figure 10. Varying $\alpha$ and $r$ in $(\alpha,r)$-$\mathsf{SPGC}$
Varying $r$ in $\mathsf{SPGC}$. Fixing $\alpha=0.25$, we vary $r$ from $1$ to $3$ over Arxiv and report the result in Fig. 10. (1) As $r$ is varied from $1$ to $3$, the inference speed-up achieved by $(0.25,r)$-$\mathsf{SPGC}$ for all $\mathsf{GNNs}$ classes notably increased. Indeed, larger $r$ allows $(\alpha,r)$-$\mathsf{SPGC}$ to find and merge more node pairs with equivalent embeddings, which may not be direct neighbors of another pair in the $(\alpha,1)$-relation. (2) As $r$ increases, the inference accuracy for all $\mathsf{GNNs}$ classes slightly drops, and all within a small range of $0.02$. This demonstrates that $(\alpha,r)$-$\mathsf{SPGC}$ is capable of preserving inference accuracy while increasing $r$ in trading for larger speed up.
Exp-3: Efficiency & Scalability. We next evaluate the compression and inference costs of $\mathsf{SPGC}$ and its variants.
Average degree and “small-world” effect v.s. Inference Speed-up. We simulate WS-MAG based on $(K,P)$ Watts-Strogatz algorithm (Watts and Strogatz, 1998) with fixed $|V|=2M$ from MAG240M dataset. $K$ and $P$ represent average degree and re-wiring probability respectively. As $P$ goes up, the less “small-world” (i.e., more random) the graphs become.
Fixing $P=0.6$, we vary the average degree from 2 to 4. Fig. 11(a) shows that the inference speed-up for different $\mathsf{GNNs}$ types. (1) As the average degree of the graph increases from $2$ to $4$, inference speed-up achieved by $\mathsf{SPGC}$ increases approximately linearly from 3.0×3.0$\times$3.0 ×-4.0×4.0$\times$4.0 × to 5.0×5.0$\times$5.0 ×-7.0×7.0$\times$7.0 ×. (2) Inference speed-up achieved by $\mathsf{DSpar}$ remains relatively stable (1.6×1.6$\times$1.6 ×-1.8×1.8$\times$1.8 ×) and smaller than $\mathsf{SPGC}$. As the average degree goes up, $\mathsf{SPGC}$ may takes the advantage of higher density that makes nodes more likely to be merged, resulting in a smaller $G_{c}$ and greater speed-up.
Next, fixing $|G|=8M$, we increase $P$ from 0.2 to 0.8. We have the following observations. (1) As $P$ increases, the inference speed-up achieved by $\mathsf{SPGC}$ on three $\mathsf{GNNs}$ exhibits a slight drop, albeit still much larger than the speed-up by $\mathsf{DSpar}$. (2) Inference speed-ups achieved by $\mathsf{DSpar}$ on three $\mathsf{GNNs}$ stay flat without even breaking 2×2$\times$2 × speed-up. This indicates that $\mathsf{SPGC}$ may perform better to compress “small-world” graphs, which are common in real-world networks such as social networks (Bassett and Bullmore, 2017; Watts and Strogatz, 1998).
(a) $|G|$ v.s. Speed-up
(b) $P$ v.s. Speed-up
(c) $|G|$ v.s. Compression Cost
(d) Compression Cost
Figure 11. Compression: Scalability and Efficiency.
Impact of $|G|$. We increased the size of the simulated graph $|G|$ from $6M$ to $600M$, spanning two orders of magnitude. As shown in Fig. 11(c), (1) $\mathsf{SPGC}$, $\mathsf{ASPGC}$, and DPP all scale well with $|G|$, which is consistent with the compression cost presented in Table. 1. (2) $(\alpha,r)$-$\mathsf{SPGC}$ has relatively higher and more sensitive compression cost as $|G|$ increases, as $(\alpha,r)$-$\mathsf{SPGC}$ incurs more time to refine the $(\alpha,r)$-relation, yet achieves better compression ratio and in turn, more reduction of the inference cost on compressed graph $G_{c}$.
Compression cost: real world datasets. Next, we compare the one-time compression cost induced by the $\mathsf{SPGC}$ or its variants with the cost of $\mathsf{DSpar}$ as shown in Fig. 11(d). Note that the compression cost of $\mathsf{SPGC}$ and its variants include the time cost of DPP. We have the following discoveries. (1) $\mathsf{SPGC}$, $(\alpha,r)$-$\mathsf{SPGC}$ and $\mathsf{ASPGC}$ outperforms $\mathsf{DSpar}$ on all the real-world datasets. In particular, $\mathsf{SPGC}$ are 95.52% and 58.45% times faster than $\mathsf{DSpar}$ over Arxiv and Products, respectively. For Yelp, $\mathsf{DSpar}$ does not run to completion after 1,000 seconds. (2) For $\mathsf{SPGC}$ and its variants, $\mathsf{ASPGC}$ outperforms $\mathsf{SPGC}$ and $(\alpha,r)$-$\mathsf{SPGC}$ by 12.71% and 68.34% on average. This is because $\mathsf{ASPGC}$ exploits data locality of $\mathsf{GNN}$ inference and perform compression up to certain hops of the anchored nodes
(a) Yelp
(b) Products
Figure 12. Total Cost: Varying Size of Queries
Total cost: Varying query workloads. We compare the total inference costs of $\mathsf{SPGC}$, $(\alpha,r)$-$\mathsf{SPGC}$, $\mathsf{ASPGC}$ to the total cost induced by the original $G$. For $\mathsf{SPGC}$ and its variants, the total cost is the sum of “one-time” compression cost and the inference cost for all inference queries; for “Original”, it refers to the inference time on the original graph. Varying the number of inference queries from $6$ to $18$ (each with fixed number of test node $|V_{T}|=0.05|V|$), Fig. 12 reports the total cost of a $3$-layers $\mathsf{GCN}$ on Yelp and Products.
[…]
(2) Among $\mathsf{SPGC}$ variants, for queries with given anchored nodes, $\mathsf{ASPGC}$ performs best in improving the inference efficiency. We also observe that $(\alpha,r)$-$\mathsf{SPGC}$ is more sensitive to different datasets, compared with $\mathsf{SPGC}$ and $\mathsf{ASPGC}$. Indeed, $\mathsf{SPGC}$ enforces rigidly embedding equivalence, while $\mathsf{ASPGC}$ benefits most from data locality from anchored node set of fixed size, hence both are less sensitive to datasets. $(\alpha,r)$-$\mathsf{SPGC}$ on the other hand is most adaptive for tunable trade-off between inference speed up and compression ratio, depending on the heterogeneity of node embeddings.
(a) Varying $\mathsf{GNN}$ Layer Number
(b) Varying $\mathsf{GNN}$ class
Figure 13. Total Cost: Varying $\mathsf{GNN}$ Layers and Class.
Total cost: varying $\mathsf{GNN}$ complexity. We next investigate the impact of $\mathsf{GNN}$ model complexity, in terms of number of layers, and types. Using $\mathsf{GCNs}$, we varying the number of layers $L$ from $2$ to $4$ (We denote ${\mathsf{GCN}}_{2}$ as the $2$-layers $\mathsf{GCN}$; similarly for others), and a query workload of $3$ different inference queries with the same size $|V_{T}|$. Fig. 13(a) tells us that as the number of layers increase, the inference cost grows as expected. $\mathsf{ASPGC}$ outperforms $\mathsf{SPGC}$ and $(0.5,1)$-$\mathsf{SPGC}$ consistently for all layers in compression efficiency.
Fixing query workload size as $3$, and the number of layers as $3$ for all $\mathsf{GNN}$ types, we report the total cost over $\mathsf{GCN}$, $\mathsf{GAT}$ and $\mathsf{GraphSAGE}$. As shown in Fig. 13(b), all three types of $\mathsf{GNNs}$ consistently and significantly benefit from $\mathsf{SPGC}$, with an improvement of inference time by a factor of $2.21$, $2.36$ and $2.11$, respectively; similar for $(\alpha,r)$-$\mathsf{SPGC}$ and $\mathsf{ASPGC}$. The one-time compression costs are consistent with our prior observations in Fig. 13(a). In general, ${\mathsf{GAT}}$ may benefit most from inference-friendly compression. A possible reason is that the memoization effectively cached its additional edge weights as part of the auxiliary information, which are a source of overhead for inference over $G$.
We have also investigated (1) how memoization structure ${\mathcal{T}}$, and neighbor recovery affect the inference accuracy and speed-up achieved by $\mathsf{SPGC}$; (2) the impact of aggregation method and improvement of model training cost. Due to limited space, we present additional tests with details in (ful, 2025).
8. Conclusion
We have proposed $\mathsf{IFGC}$, an inference friendly graph compression scheme to generate compressed graphs that can be directly processed by $\mathsf{GNN}$ inference process to obtain original inference output. We have introduced a sufficient condition, and introduced three practical specifications of $\mathsf{IFGC}$, $\mathsf{SPGC}$, for inference without decompression, and configurable $(\alpha,r)$-compression, to achieve better compression ratio and reduction of inference cost, and anchored $\mathsf{SPGC}$, that preserves inference results for specified node set. Our theoretical analysis and experimental study have verified that $\mathsf{IFGC}$-based approaches can significantly accelerate the inference on real-world and large graphs with small loss on inference accuracy. A future topic is to extend our compression scheme to accelerate $\mathsf{GNN}$ training. Another topic is to develop parallel compression scheme for large graphs.